3. Great Britain and the United States produce cheddar cheese and blue cheese. Current domestic prices per pound for each type of cheese are given in the following table: Great Britain 210 £20 States Blue cheese $50 Suppose the exchange rate is £1-$1 If the price ratios within each country relect resource use, (1) has a comparative advantage in the production of cheddar cheese. (2)- has a comparative advantage in the producton of blue cheese. Assume there are no other trading partners and that the only motive for holding foreign currency is to buy foreign goods. Explain whether the current exchange rate will lead to trade flows in both directions between the two countries OA. Cheddar cheese and blue cheese will be purchased in the United States and sold in O B. Only cheddar cheese produced in the United States will be traded for blue cheese produced in O C. Cheddar cheese and blue cheese will be purchased in Great Britain and sold in O D. Only cheddar cheese produced in Great Britain will be traded for blue cheese produced in Great Britain. Great Britain. the United States the United States Trade will not occur E. What adjustments might you expect in the exchange rate? Relative to the British pound, the US. dollar will (3) After the exchange rate has adjusted, (4) produced in Great Britain wil be consumed in the United States and (5) produced in the United States will be consumed in Great Britain. (1) Great Britain(2) O United States (3) O become weaker (4) O O Great Britain blue cheese cheddar cheese (5) O cheddar cheese Oblue cheese O United States O remain unchanged O become strongerPlease help me with Economics Homework?
Answer #1

cheddar cheese price in Britain = 10 pounds . as exchange rate is 1 pound = 1$, so price of cheddar cheese = $10

similarly price of blue cheese = $20

Great Britain , cheddar cheese/ blue cheese = 1/2 = 0.5

United States , cheddar cheese/ blue cheese = 35/50 = 0.7, since 0.5< 0.7

1)so Great Britain has comparative advantage in cheddar cheese

Great Britain , blue cheese/cheddar cheese= 20/10 = 2

United States , blue cheese/cheddar cheese= 50/35 = 1.42 , since 1.42 < 2 so 2)United States has comparative advantage in blue  cheese .

at current exchange rate , both blue and cheddar cheese are cheaper in Great Britain , so it will be purchased in great britain and sold,in United States i.e option C is answer

so relative to pound , dollar will 3)become weaker as it supply > demand so dollar will depreciate

after adjustment , 4)cheddar. cheese produced in Britain will be consumed in United States( as Britain has. a comparative advantage in this ) and 5)blue cheese produced in United States will be consumed in Britain ( as united States  has. a comparative advantage in this)